The probability of a snowstorm on thanksgiving is .20 and the probability of a snowstrom on christmas is .90. Assuming that these two events are independent, what is the probability of a snowstorm on either thanksgiving on christmas?
Answer #1

Ans

P(snowstorm on either thanksgiving on Christmas) = 1- P(nither) = 1 - (0.8*0.1) = 0.92